OpenStudy (anonymous):

Find the slope of the tangent line to the curve sqrt(4x+1y)+sqrt(4xy)=13.7 at the point (6,3)

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

the square roots really throw me off, so idk :/

OpenStudy (anonymous):

ick does it really say \[\sqrt{4x+y}+\sqrt{4xy}=13.7\]?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

lol your math teacher must hate you

OpenStudy (anonymous):

yessss lol

OpenStudy (anonymous):

before i write all the muck we gotta write, lets at least make \(\sqrt{4xy}=2\sqrt{xy}\) okay?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

and also recall that the derivative of \(\sqrt x\) is \(\frac{1}{2\sqrt{x}}\)

OpenStudy (anonymous):

now just grind it out \[\frac{4+y'}{2\sqrt{4x+y}}+\frac{y+xy'}{\sqrt{xy}}=0\]

OpenStudy (anonymous):

probably went with too few steps any idea how i got that?

OpenStudy (anonymous):

okay so the derivative of \[\sqrt{4x+y}\]=4+y' then divided by 2\[\sqrt{4x+y}\]

OpenStudy (anonymous):

right

OpenStudy (anonymous):

the second part im kind of confused

OpenStudy (anonymous):

ok well the twos cancelled for one

OpenStudy (anonymous):

we pulled out the two, and that goes with the two in the denominator

OpenStudy (anonymous):

then for the derivative of \(xy\) we need the product rule

OpenStudy (anonymous):

oh okay, got it

OpenStudy (anonymous):

good so the derivative of \(xy\) wrt x is \(y+xy'\) via the product rule, and that goes in the numerator

OpenStudy (anonymous):

now replace x by 6,y by 3 and solve for \(y'\)

OpenStudy (anonymous):

so, \[\frac{ 4+y' }{ 6\sqrt{3} }+\frac{ 3+6y' }{ 3\sqrt{2} }=0\]?

OpenStudy (anonymous):

would you multiply both denominators?

OpenStudy (anonymous):

@jim_thompson5910 ?

jimthompson5910 (jim_thompson5910):

yeah you could multiply both sides by the LCD \[\Large 6\sqrt{6}\] to clear out the fractions

OpenStudy (freckles):

sqrt(4x+y)+sqrt(4xy) is approximately 13.7 at (x=6,y=3)

OpenStudy (anonymous):

i got y'=-1, it says its wrong

OpenStudy (freckles):

so what I'm saying is this point is not totally on any tangent line to the curve

OpenStudy (freckles):

it would be really close but not on

OpenStudy (anonymous):

now im confused lol

OpenStudy (freckles):

\[\frac{4+y'}{2\sqrt{4x+y}}+\frac{y+xy'}{\sqrt{xy}}=0 \\ \frac{4+y'}{2 \sqrt{4(6)+3}}+\frac{3+6y'}{\sqrt{6(3)}}=0 \\ \frac{4+y'}{2 \sqrt{27}}+\frac{3+6y'}{\sqrt{18}}=0 \\ \frac{4+y'}{2 \sqrt{9 \cdot 3}}+\frac{3+6y'}{\sqrt{9 \cdot 2}}=0 \\ \frac{4+y'}{6 \sqrt{3}}+\frac{3+6y'}{3 \sqrt{2}}=0 \\ \text{ multiply both sides by } 6 \sqrt{6} \text{ as suggested by jim } \\ 6 \sqrt{6} \frac{4+y'}{6 \sqrt{3}}+6 \sqrt{6} \frac{3+6y'}{3 \sqrt{2}}=0 \\ \sqrt{2} (4+y')+2 \sqrt{3}(3+6y')=0 \\ 4 \sqrt{2} +\sqrt{2}y' +6 \sqrt{3}+12 \sqrt{3}y'=0\] and you think this will end up being y'=-1? but anyways I think this question itself is a trap because as I said (6,3) is not even on the curve it is near but not on

OpenStudy (anonymous):

yeah i got y'=-1.19112, what do you get?

OpenStudy (anonymous):

@freckles

OpenStudy (freckles):

\[\sqrt{2}y'+12 \sqrt{3}y'=-4 \sqrt{2}-6 \sqrt{3} \\ (\sqrt{2}+12 \sqrt{3})y'=-4 \sqrt{2}-6 \sqrt{3} \\ y'=\frac{-4 \sqrt{2}-6 \sqrt{3}}{\sqrt{2}+12 \sqrt{3}}\] but as i said before finding y' here shouldn't matter because the point (6,3) is not even on the given curve so you shouldn't be able to find the tangent to the given curve at that point ...

OpenStudy (anonymous):

i guess my professor did a mistake then, i emailed him already

OpenStudy (anonymous):

thanks anyways